Grade 12th passMechanics

Train consisting of 30 Wagon each of 15 tons and is pulled by an engine of power 90 kilowatts if the frictional force and air resistance force offered per ton of the train is 12 Newtons. The maximum velocity with which train can travel is

AKOJU SAI KUMAR
8 Years agoGrade 12th pass

1 Answer

a dog
3 Years ago
Total mass = 30 × 15 = 450 tonnes 
Power = 90 × 10watts = F × V
 Given, frictional force per unit = 12 N 
Therefore, F / 450 =12 ,  F = 450 × 12 , 
Maximum velocity = 90000 / (450×12) = 16.66 m/s